# A Calculus of Inheritance

### 10 Mar 2026 v6

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. By unifying classes, methods, objects, and properties under a single record abstraction, the calculus models inheritance simply as set union. Consequently, composition is inherently commutative, idempotent, and associative, structurally eliminating the multiple-inheritance linearization problem. Its semantics is first-order, denotational, and fixpoint-computable. These three properties extend to the lazy $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for a sublanguage of inheritance-calculus, except that recursive $\lambda$-programs that diverge under $\beta$-reduction may converge under $\mathrm{lfp}(T_{P})$. Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
@@
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of this table. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra. We challenge this assumption. Configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation. The nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise. Inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one. Prior systems such as Scala and the NixOS module system assume single-object resolution and reject the multi-path situation. Self-referential records require no fixed-point combinator in the object language. The mixin tree is the call stack, and qualified this resolution is a first-order set-theoretic query over paths. We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism. Church-encoded values automatically form trie structures, and operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$—the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5). Furthermore, the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015) (Section Function color blindness: ).
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the lazy $\lambda$-calculus cannot macro-express the inheritance constructors. Inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing that inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) to serve as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) to serve as the foundation of declarative programming.
All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example. [^1] MIXINv2 separates symbol resolution (which resolves lexical references and qualified-this references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense. Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
Contributions
  • We present inheritance-calculus, a minimal computational model for declarative programming that contains three constructs (record, definition, inheritance).
  • The mixin is a single abstraction that unifies class, method, object, and property. Method-level overriding is subsumed by deep merge of mixin trees; inheritance is commutative, idempotent, and associative, and the linearization problem in OOP does not arise.
  • The semantics is observational: the observer drives computation by querying paths in a lazily constructed tree. The six semantic equations are first-order [^2] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies; the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions; and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates., denotational, and fixpoint-computable.
  • The mixin tree is the call stack. Self-reference resolves to multiple objects. Since inheritance is idempotent, all objects contribute equally.
  • The $\lambda$-calculus macro-expresses into inheritance-calculus in five local rules (Section 4).
  • The translation gives the lazy $\lambda$-calculus a first-order semantics (Section 4.1): the prior higher-order semantic frameworks are shown to be faithful to a sublanguage of inheritance-calculus’s first-order equations. Appendix E reformulates this semantics directly on ANF. As a corollary, some $\lambda$-terms that diverge in the $\lambda$-calculus converge under $\mathrm{lfp}(T_{P})$ after translation (Section 4; Remark B.5).
  • Inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
  • We observe further properties: immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
Section 5 demonstrates these properties through a case study of Nat arithmetic.
@@ 2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a nonnegative integer (path length).
@@
$\mid\quad$ $[\ell_{\mathrm{up}}.\mathbf{this}.\mid\uparrow^{\!n}.]\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ $\text{(inheritance, }k\geq 0\text{ with qualifier,}$
$k\geq 1\text{ without)}$
@@
The $\mapsto$ in $\ell\mapsto e$ defines a property. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
@@
A new scope level is created at each $\{\ldots\}$ record literal. The scope contains all properties of that mixin, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single mixin whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
@@
  • $\mathrm{Foo}.\mathbf{this}.\ell$ (equivalently, $\uparrow^{\!n}.\ell$) projects property $\ell$ from $\mathrm{Foo}$.
@@ 3. Mixin Trees
In object-oriented programming, a mixin (Bracha and Cook, 1990) is a fragment of behavior that can be composed onto a class via inheritance. Inheritance-calculus has a single abstraction, the deep-mergeable mixin, [^3] Where no ambiguity arises, we write simply “mixin.” The distinction from the Bracha–Cook mixin (Bracha and Cook, 1990) is detailed in Section 7. which can simultaneously serve as field, method, class, and object, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees. We call the resulting structure a mixin tree. Its semantics is purely observational: given an inheritance-calculus program, one queries whether a label $\ell$ is a property at a given position $p$.
This section defines the semantics via six mutually recursive first-order functions on paths. The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and self-reference.
@@ 3.1. Definitions
  • $\mathtt{defines}(p)$: the set of labels that have subtrees at $p$.
@@
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the $\mathtt{overrides}$ of $p$ itself, plus the $\mathtt{overrides}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified this resolution, i.e. the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
@@
When multiple routes exist due to multi-path inheritance, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified this resolution
@@
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates. [^4] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target, and no properties are inherited. This is the most self-consistent semantics for the untyped calculus. MIXINv2 (Yang, 2026) is stricter: its resolver rejects references whose frontier becomes empty, reporting a resolution error at compile time.
@@ 3.2. Recursive Evaluation
The six equations (1)–(6) define $\mathtt{properties}(p)$ by mutual recursion: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these functions. The functions are the interpreter—the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand. Because the functions are pure and order-independent, the reflexive-transitive closure $\mathtt{bases}^{*}$ in (2) may be explored breadth-first or depth-first without affecting the result; paths may be interned so that structural equality reduces to pointer equality; and the six functions form a memoizable dynamic-programming recurrence keyed by paths.
Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996): each query is memoised on first encounter, and re-entrant calls (cycles in the dependency graph) return the current accumulator rather than diverging. Denotationally, the six equations define a monotone immediate consequence operator $T_{P}$ (van Emden and Kowalski, 1976) whose right-hand sides use only $\cup$, $\in$, equality, and positive set comprehension, with no negation, set difference, or complement. By the Knaster–Tarski theorem (Tarski, 1955), $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$, which serves as the denotational semantics. Appendix A proves that tabled evaluation never returns a wrong answer and analyses termination.

### 16 Mar 2026 v7

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. By unifying modules, classes, objects, methods, fields, and locals under a single record abstraction, the calculus models inheritance simply as set union. Consequently, composition is inherently commutative, idempotent, and associative, structurally eliminating the multiple-inheritance linearization problem. Its semantics is first-order (van Emden and Kowalski, 1976; Reynolds, 1972; Aczel, 1977), denotational, and computable by tabling (Tamaki and Sato, 1986), even for cyclic inheritance hierarchies. These three properties extend to the $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for the first-iteration approximation of a sublanguage of inheritance-calculus. As a corollary, this establishes a convergence hierarchy $\text{eager}\subsetneq\text{lazy}\subsetneq\text{fixpoint}$ among $\lambda$-calculi sharing the same $\lambda$-syntax.
Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); $\mathtt{this}$ resolves to multiple targets; and programs are immune to nonextensibility in the sense of the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, ), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, ) allows independently authored modules to extend the system configuration without modifying existing modules; in practice, the ecosystem has not been troubled by the multi-dimensional extension difficulty analogous to the Expression Problem (Wadler, 1998). [^1] As a rough proxy for ecosystem scale: according to Repology (Repology Contributors, 2025), nixpkgs contains over 113,000 packages, nearly three times the count of Debian or Ubuntu, with a comparable number of contributors. Not every package exercises the module system’s extensibility, but the same caveat applies to other distributions. The mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory captures the extensibility that this mechanism exhibits.
The $\lambda$-calculus serves as the foundational computational model for functional programming. No analogous calculus exists for declarative programming. This gap matters because the two paradigms differ in fundamental ways. Configuration languages are inherently declarative: their values are immutable and they have no first-class functions. Note that a function body is closed: its internal structure cannot be extended without modifying the definition, whereas the configuration languages above all compose by extending existing structure from outside. At first glance, Turing completeness appears incompatible with these constraints. The three classical models of computation each violate at least one of them:
@@
The Turing Machine and RAM Machine require mutable state; the $\lambda$-calculus requires first-class functions. A computational model that is Turing complete, immutable, and function-free would sit in the empty cell of Table 1. The NixOS module system already suggests that such a model exists: its inheritance-based composition over recursive records, without explicit functions, is expressive enough to configure entire operating systems.
@@
Most configuration languages operate on finite, well-founded structures (initial algebras in the sense of universal algebra). Nix is an exception: although designed specifically as a configuration language for NixOS (Dolstra, 2006), it is a complete programming language whose values are lazily observable and possibly infinite. The nixpkgs package collection (over 100,000 packages) behaves as a lazily observed structure, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, definitions, and inheritance) and no functions or let-bindings. [^2] The executable MIXINv2 implementation (Yang, 2026) contains every example in this paper.
@@
In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, several phenomena arose from off-label use of the adopted design patterns. Section 3 shows how the three constructs give rise to a single abstraction that simultaneously serves as module, class, object, method, field, and local binding. The linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise: inheritance is inherently commutative, idempotent, and associative, and $\mathtt{this}$ naturally resolves to multiple targets rather than one, where prior systems such as Scala and the NixOS module system reject the multi-target situation. Section 5 traces how boolean logic and arithmetic, implemented as separate files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism, and how ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976). Section 6.1 discusses these and further emergent phenomena, including function color blindness (Nystrom, 2015).
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Section 3 defines the semantics via six mutually recursive first-order functions on paths and sets of paths, with no function spaces, closures, or higher-order constructs; the functions are computable by tabling (Tamaki and Sato, 1986). Section 4 shows that these functions extend to the $\lambda$-calculus, embedding it. Remark B.5 derives a convergence hierarchy among eager, lazy, and fixpoint $\lambda$-calculi sharing the same syntax as a corollary. Section 4.2 establishes that the lazy $\lambda$-calculus cannot macro-express the inheritance constructors, so inheritance-calculus is strictly more expressive in Felleisen’s sense (Felleisen, 1991).
@@ 2. Syntax
Let $e$ denote an expression, $\ell$ a label serving as a property name, and $k$ a nonnegative integer serving as the length of a downward projection.
@@
$\mid\quad$ $(\ell_{\mathrm{up}}.\mathbf{this}\mid\uparrow^{\!n}).\;\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ (inheritance: qualifier + downward path)
$\mid\quad$ $\ell_{\mathrm{up}}.\mathbf{this}\mid\uparrow^{\!n}$ (inheritance: qualifier only)
$\mid\quad$ $\ell_{\mathrm{down},1}.\ell_{\mathrm{down},2}\ldots\ell_{\mathrm{down},k}$ (inheritance: downward path only)
@@
The $\mapsto$ in $\ell\mapsto e$ defines a property named $\ell$ with value $e$. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; no scalar values are needed in the examples.
@@
A new scope level is created at each $\{\ldots\}$ record literal. [^3] We use record for the syntactic construct, scope for the name-resolution context it creates, and mixin for its role as a composable unit; in inheritance-calculus these are three views of the same entity. The scope contains all properties of that mixin, including those inherited via inheritance sources. For example, $\{a\mapsto\{\},\;b\mapsto\{\}\}$ is a single mixin whose scope contains both $a$ and $b$; so is $\{r,\;a\mapsto\{\}\}$ where $r$ is a reference to a record containing $b$. Inheritance sources do not create additional scope levels.
@@
  • $\mathrm{Foo}.\mathbf{this}.\ell$, or equivalently, $\uparrow^{\!n}.\ell$, projects property $\ell$ from $\mathrm{Foo}$.
@@ 3. Mixin Trees
In object-oriented programming, a mixin (Bracha and Cook, 1990) is a fragment of behavior that can be composed onto a class via inheritance. Inheritance-calculus has a single abstraction, the deep-mergeable mixin, [^4] Where no ambiguity arises, we write simply “mixin.” The distinction from the Bracha–Cook mixin (Bracha and Cook, 1990) is detailed in Section 7. which can simultaneously serve as module, class, object, method, field, and local binding, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees. We call the resulting structure a mixin tree. Its semantics is purely observational: given an inheritance-calculus program, one queries whether a label $\ell$ is a property at a given position $p$.
This section defines the semantics via six mutually recursive first-order [^5] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies (van Emden and Kowalski, 1976); the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions (Reynolds, 1972); and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates. functions on paths. The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and qualified $\mathtt{this}$ resolution.
@@ 3.1. Definitions
  • $\mathtt{defines}(p)$: the set of labels $\ell$ for which $p$ contains a definition $\ell\mapsto e$.
@@
Intuitively, $\mathtt{supers}(p)$ collects every path that $p$ inherits from: the $\mathtt{overrides}$ of $p$ itself, plus the $\mathtt{overrides}$ of each direct base $\mathtt{bases}$ of $p$, and so on transitively. Each result is paired with the inheritance-site context $\mathtt{init}(p_{\mathrm{base}})$ through which it is reached. This provenance is needed by qualified $\mathtt{this}$ resolution, that is, the $\mathtt{this}$ function defined below, to map a definition-site mixin back to the inheritance-site paths that incorporate it:
@@
When multiple routes exist due to multi-target $\mathtt{this}$ resolution, they may yield different target paths; all are collected. We now define $\mathtt{this}$.
Qualified $\mathtt{this}$ resolution
@@
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates. [^6] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target; no properties are inherited. This is the most self-consistent semantics for the untyped calculus. MIXINv2 (Yang, 2026) is stricter: its resolver rejects references whose frontier becomes empty, reporting a resolution error at compile time.
@@ 3.2. Recursive Evaluation
The six equations (1)–(6) define $\mathtt{properties}(p)$ by mutual recursion: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these functions. The functions are the interpreter: the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand. Because the functions are pure and order-independent, the reflexive-transitive closure $\mathtt{bases}^{*}$ in (2) may be explored breadth-first or depth-first without affecting the result; paths may be interned so that structural equality reduces to pointer equality; and the six functions form a memoizable dynamic-programming recurrence keyed by paths.
Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986): each query is memoised on first encounter, and re-entrant calls, which correspond to cycles in the dependency graph, return the current accumulator rather than diverging. Denotationally, the six equations define a monotone immediate consequence operator $T_{P}$ (van Emden and Kowalski, 1976) whose right-hand sides use only $\cup$, $\in$, equality, and positive set comprehension, with no negation, set difference, or complement. By the Knaster–Tarski theorem (Tarski, 1955), $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$, which serves as the denotational semantics. This denotational characterization formalizes the observational semantics introduced above: the observer-driven queries are computed by the same fixed-point equations. Whether evaluation terminates depends on the program. Recall that $\mathtt{supers}$ (equation 2) takes the reflexive-transitive closure of $\mathtt{bases}$ (equation 4); the resulting graph is the inheritance hierarchy. Since each inheritance source occupies a distinct node in the mixin tree, this hierarchy is directly readable from the tree structure. Programs whose inheritance hierarchy is acyclic terminate in one pass. When the hierarchy contains cycles, evaluation terminates when each query has a finite answer domain: the ascending chain condition on the finite lattice guarantees convergence (Bancilhon and Ramakrishnan, 1986), as in Datalog (Appendix F). When the hierarchy is infinite, evaluation may diverge; this is inherent to Turing completeness. Appendix A proves that tabled evaluation never returns a wrong answer and formalizes these termination conditions.